(1 point) In this problem you will solve the differential equation (+7)y+11xy - y=0. x for the differential equation will
Answer #1

R-PIFE (y t-2n y- n=2 う Pat abone Values C0 Sn(nプ)enxと 十11x5nenx -2. Cxta) - D - ax= N3 n= ハ=D MullApiyiry the sum cofficent00 00 メt 2ain-) Cnx +4 Cg +432メ+2 (nt>) LnナDnt n-2 の=2 - o - Co-CIX Gx duelly Combuning ハ-Q From conslaut eun n eveh Know dh10 Saso 70 8.2, 520 tne STution s Co t Cix+ Gxxlu Cot Gx tloyx-Co3304x C 82,82 2 176 82320